OpenStudy (anonymous):

if you deposit $1,000 in an account that pays 6% annual interest compounded continuously, what will the balance be after 5 years

OpenStudy (anonymous):

A = P (1 + R/100)^n

OpenStudy (anonymous):

A = 1000 (1 + 6/100)^5

OpenStudy (anonymous):

A = 1000(106/100)^5 = 1000 (53/50)^5 = 1000(418195493/312500000) 1000 * 418195493 418195493 A = ------------------- = ------------ = 1338.23 approx. 312500000 312500

OpenStudy (anonymous):

thats what i got but that answer choices are A 1,932.06 B 1,349.86 C 29,803.91 D 1,232.06

OpenStudy (anonymous):

Well,..... there can always be a printing error...☺ or the nearest one is B. 1349.86